OpenStudy (cutiecomittee123):

Simplify (-10-sqrt125)/25

OpenStudy (misty1212):

HI!!

OpenStudy (cutiecomittee123):

Hello can you help please!

OpenStudy (misty1212):

\[125=25\times 3\] so \[\sqrt{125}=\sqrt{25\times 3}=\sqrt{25}\sqrt3=5\sqrt3\]

OpenStudy (cutiecomittee123):

Don't you mean 25 x 5?

OpenStudy (misty1212):

so you have \[\frac{-10-5\sqrt5}{5}\] as a start divide each by 5

OpenStudy (misty1212):

oops yeah that is what i meant sorry last box is correct though

OpenStudy (cutiecomittee123):

Also what happened to the 25? How did it turn into a 5

OpenStudy (misty1212):

\[\frac{-10-5\sqrt5}{5}\] each gets divided by \(5\)

OpenStudy (misty1212):

it turned in to a five because you were taking the square root of 25 \[\large \sqrt{125}=\sqrt{25\times 5}=\color{red}{\sqrt{25}}\sqrt5=\color{red}5\sqrt5\]

OpenStudy (cutiecomittee123):

If each was divided by 5 then wouldn't it be (-2 - sqrt5)/5

OpenStudy (misty1212):

no

OpenStudy (misty1212):

when you divide by 5, there is no more denominator

OpenStudy (misty1212):

\[-\frac{10}{5}=-2\\ -\frac{5\sqrt5}{5}=-\sqrt5\]

OpenStudy (misty1212):

final answer \[-2-\sqrt5\] the \(5\) in the denominator is gone

OpenStudy (hocuspucus292333393):

can u help me misty please

OpenStudy (cutiecomittee123):

Okay I see

OpenStudy (misty1212):

oh wait hold on i made a mistake sorry

OpenStudy (misty1212):

the denominator is 25, not 5, you were right to begin with

OpenStudy (misty1212):

your answer was correct, only a five cancels so you were correct \[\frac{-2-\sqrt5}{5}\]

OpenStudy (alirayyan):

yes , this one is correct :) @misty1212